DIRECT OBSERVATION OF DNA REPAIR IN PHOTOLYASE

\underline{YA-TING KAO}, CHAITANYA SAXENA, LIJUAN WANG and DONGPING ZHONG, Department of Physics, Chemistry and Biochemistry, OSU Biophysics, Chemical Physics and Biochemistry Programs, The Ohio State University, Columbus, Ohio 43210.

Photolyase splits cyclobutane ring of pyrimidine dimer (CPD) in DNA in a light (350-500nm) driven reaction and thus reverses the harmful effects of far-UV (200-300nm). Through resonance energy transfer from photoantenna MTHF or directly by absorption of visible light, reduced flavin cofactor (FADH^{−}) transfers an electron to CPD which subsequently splits into two pyrimidines. Concomitantly, an electron is transferred back to the oxidized neutral radical (FADH^{o}) to restore the active form (FADH^{−}). With femtosecond resolution we followed the functional evolution and observed forward and backward electron transfer reactions as well as the entire DNA-repair process. These dynamics occur in the picosecond time scale, raveling the ultrafast nature of DNA-repair mechanism